# Linear cost and exponentially convergent approximation of Gaussian Matérn processes

### v1

@@ 4 Numerical results
In this section, we illustrate the performance and accuracy of the proposed rational approximation method by comparing it with several alternative approaches. Specifically, we compare our method (referred to as Rational Approximation) with the state-space method of Karvonen & Särkkä (2016), the nearest-neighbor Gaussian process approximation (referred to as nnGP) of Datta et al. (2016), the random Fourier features method (referred to as Fourier) from Rahimi & Recht (2007), a principal component analysis approach (referred to as PCA) to serve as a lower bound for low-rank methods, the tapering method (referred to as Taper) of Furrer et al. (2006), and the covariance-based rational SPDE approach (referred to as FEM) from Bolin et al. (2024b).
We compare the accuracy of the proposed method with the alternative methods by carrying out three tasks. First, we measure and compare the accuracy of our method with the alternative methods in terms of the accuracy of their covariance function, and then consider the cost of Gaussian process regression and measure the quality in terms of the accuracy of the posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$. Finally, we compare the accuracy of the process approximation as a whole by computing coverage probabilities of joint confidence bands in a Gaussian process regression. In all these cases, we consider the accuracy for a fixed computational cost (as explained below) to make the comparison completely fair. Further, all methods were implemented in R, using the same methods for sparse matrices, matrix solves, and other tasks. The results were obtained using a MacBook Pro Laptop with an M3 Max processor and 128Gb of memory, without using any parallel computations to make the comparison as fair as possible.
Table 1: Asymptotic costs for different methods. For nnGP, $m$ is the number of neighbors and for the low rank methods, $m$ is the rank. For the state-space method, the “preconditioned” approximation of Karvonen & Särkkä (2016) is used in combination with our Markov approach, see Remark 3.6. All costs are in “Big O” assuming $n$ is much larger than $\alpha$ and $m$. For FEM, $N$ is the number of mesh nodes.
@@
state-space (Karvonen & Särkkä, 2016) $n(m+\lfloor\alpha\rfloor)\lceil\alpha\rceil^{4}$ $n(m+\lfloor\alpha\rfloor)\lceil\alpha\rceil^{2}$ $n(m+\lfloor\alpha\rfloor)^{3}\lceil\alpha\rceil^{3}$
@@
Fourier (Rahimi & Recht, 2007) $nm$ $nm$ $nm^{2}$
@@ 4.1 Setup for the first comparison
We first consider the case of a dataset with 5000 observations, ${\boldsymbol{\mathrm{y}}}=[y_{1},y_{2},\dotsc,y_{n}]$, distributed over the interval $[0,50]$. Each observation is generated as $y_{i}\mid u(\cdot)\sim\mathcal{N}(u(t_{i}),\sigma_{e}^{2}),$ where $t_{i}$ are the observation locations, and $u$ is a centered Gaussian process with the Matérn covariance function given by (1). We set $\sigma=1$ (as it is merely a scaling parameter), fix $\sigma_{e}=0.1$, and vary $\nu$ over the interval $(0,2.5)$. For each value of $\nu$, we choose $\kappa=\sqrt{2\nu}$, ensuring that the practical correlation range, $\rho=\nicefrac{{\sqrt{8\nu}}}{{\kappa}}$, remains fixed at 2.
This choice of range, along with the interval length, ensures numerical stability across all methods compared. Specifically, a practical correlation range of 2 was the largest range for which all methods, including the most sensitive (nnGP), remained stable. While a practical correlation range of 2 on a domain of size 50 may seem small, stability and accuracy of the predictions depend on the number of observations within the correlation range at any given point, rather than the size of the domain itself. Other ranges, and other numbers of observation and prediction locations, are investigated in the Shiny app.
The goal is to compute the posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ with elements $({\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}})_{j}=\mathbb{E}(u(p_{j})|{\boldsymbol{\mathrm{y}}})$, where $p_{j}$ are the prediction locations. To make the comparison simple, we initially choose $p_{j}=t_{j}$ evenly spaced in the interval. Because the locations are evenly spaced, we could also include the Toeplitz method of (Ling & Lysy, 2022) in the comparison. However, we do not include that here as we only consider generally applicable methods.
@@ 4.2 Calibration of the computational costs
The asymptotic costs of the different methods are summarized in Table 1. However, we calibrate the methods to have the same total runtime for assembling all matrices (the construction cost) and computing the posterior mean (the prediction cost). This ensures fair comparisons based on actual performance rather than theoretical complexity, which can be affected by the constants in the cost expression and the size of the study. Specifically, for a given set of parameters $(\kappa,\sigma,\nu)$, and a fixed value of $m$ for the proposed method, we calibrate the values of $m$ for the other methods to ensure that the total computation time is the same. The total prediction times were averaged over 100 samples to obtain the calibration. The final calibration results are shown in Table 2. We now provide some details regarding this calibration.
@@
As described in Remark 3.6, the state-space method provides an alternative Markov representation for which we could use the same computational methods as for the proposed method. Its value of $m$ was therefore chosen according to Table 1 as $m-\lfloor\alpha\rfloor$.
To minimize boundary effects of the FEM method, we extended the original domain $[0,50]$ to $[-4\rho,50+4\rho]$, where $\rho$ is the practical correlation range. This extension ensures accurate approximations of the Matérn covariance at the target locations. Because the FEM method uses the same type of rational approximation as the propose method, we fixed the value of $m$ for the FEM method to be equal to $m$ for the proposed method. The calibration was then instead performed on the finite element mesh. Specifically, a mesh with $N=kn+500-(k+3)$ locations in the extended domain used, where $500$ locations were in the extensions and $kn$ locations in the interior $[0,50]$, and the $-(k+3)$ term appears to ensure that the regular mesh contains the observation locations. These $kn$ locations were chosen equally spaced to include the observation locations and $k\in\mathbb{N}$ was calibrated to make the total computational cost match that of the proposed method for $\nu<1.5$, and for $2.51.5<\nu<$ 2.5 it was chosen as the largest values that yielded a stable prediction, as the value which matched the computational cost yielded unstable predictions. The resulting values of $N$ are shown in Table 3.
4.3 Results
@@
$\|r-\hat{r}\|_{L_{2}}^{2}\approx\frac{1}{n^{2}}\sum_{i,j=1}^{n}(r(t_{i},t_{j})-\hat{r}(t_{i},t_{j}))^{2},\quad\|r-\hat{r}\|_{L_{\infty}}\approx\max_{i,j}|r(t_{i},t_{j})-\hat{r}(t_{i},t_{j})|,$
where $r$ is the Matérn covariance (1) and $\hat{r}$ is the approximation obtained by the considered methods. We also consider the error of the actual prediction, To compute the prediction error, we first compute the true posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ and then compute the corresponding posterior mean $\hat{{\boldsymbol{\mathrm{\mu}}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ under each approximate model and the corresponding $L_{2}$ error
$\|{\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}-\hat{{\boldsymbol{\mathrm{\mu}}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}\|_{L_{2}}^{2}=\frac{1}{n}\sum_{i=1}^{n}(({\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}})_{i}-(\hat{{\boldsymbol{\mathrm{\mu}}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}})_{i})^{2}.$
@@
In Figure 1, the resulting $L_{2}$ and $L_{\infty}$ covariance errors can be seen for the different methods when $n=5000$. The left panel shows the error for our method with different choices of $m$, and one can see that the covariance error indeed decreases rapidly to zero as $m$ increases. In the remaining two panels, we show the error for $m=3$ and $m=5$ and compare these errors to the error from the competing methods, calibrated to have the same cost of prediction as described above. We can note that the competing methods are less accurate and in fact, one can observe that the nnGP method becomes slightly unstable for large values of smoothness parameter $\nu$, whereas the FEM method becomes very unstable due to the calibrated mesh being very fine. It is also noteworthy that the mesh is fine enough that when the method is stable (which mean $\nu<0.9$ for $m=3$ and $\nu<0.5$ for $m=5$), we can only see the rational approximation error, as both FEM and rational methods overlap. This shows that the FEM method is an excellent choice when stable, but that the rational is better because it is more stable and does not require choosing a mesh or a boundary extension. The PCA method is more accurate for large values of $\nu$, but one should recall that it is a theoretical lower bound for low rank methods, and not a practically competitive method, as it has an $\mathcal{O}(n^{3})$ construction cost unless the eigenvectors are known explicitly. Any practically useful low rank method, such as the process convolution approach (Higdon, 2002) or fixed rank kriging (Cressie & Johannesson, 2008), will have larger errors, as can clearly be seen when considering the errors of the random Fourier features method. Finally, one should note that for values $\nu=0.5,1.5,2.5$ in the considered interval $(0,3)$, the proposed method is exact.
@@
Figure 2: Prediction errors measured in mean square difference on interval $I=[0,50]$ for $\rho=2$, different values of $\nu$ and different orders of approximation.
The prediction errors shown in Figure 2 further demonstrate the advantages of the rational approximation method over the other approaches. As the approximation order $m$ increases, the mean squared error decreases consistently and rapidly to zero. For $m=3$ we can see that nnGP outperforms the rational method, by a small margin, for $\nu<1$ and the two methods have similar performance for $\nu\geq 1$. For $\nu<0.5$, the rational and FEM methods overlap. Moving to $m=5$, we can see that nnGP outperforms the rational method for small values of the smoothness parameter $\nu$, but the rational method outperforms nnGP for $\nu>0.8$. In this case, there is also an overlap between FEM and rational methods for $\nu<0.5$. It is noteworthy that the quality of the rational approximation we have available deteriorates as $\nu$ approaches zero, which suggests us that it might be possible to improve the approximation for $\nu<0.5$ by obtaining better coefficients for rational approximations in such cases. Finally, FEM provides very good accuracy, comparable to the rational and nnGP methods, whereas PCA provides the best results for large values of $\nu$, however, we recall that for one to be able to use PCA with this cost, the eigenvectors must be known.

### v2

@@ 4 Numerical results
In this section, we illustrate the performance and accuracy of the proposed rational approximation method by comparing it with several alternative approaches. Specifically, we compare our method (referred to as Rational Approximation) with the state-space method of Karvonen and Särkkä (2016), the nearest-neighbor Gaussian process approximation (referred to as nnGP) of Datta et al. (2016), the random Fourier features method (referred to as Fourier) from Rahimi and Recht (2007), a principal component analysis approach (referred to as PCA) to serve as a lower bound for low-rank methods, the tapering method (referred to as Taper) of Furrer et al. (2006), and the covariance-based rational SPDE approach (referred to as FEM) from Bolin et al. (2024b).
We compare the accuracy of the proposed method with the alternative methods by carrying out three tasks. First, we measure and compare the accuracy in terms of the accuracy of the covariance function, and then consider the cost of Gaussian process regression and measure the quality in terms of the accuracy of the posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ and posterior standard deviation ${\boldsymbol{\mathrm{\sigma}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$. Finally, we compare the accuracy of the process approximation as a whole by computing Kullback–Leibler (KL) divergences and coverage probabilities of joint confidence bands in a Gaussian process regression. In all these cases, we consider the accuracy for a fixed computational cost (as explained below) to make the comparison completely fair. Further, all methods were implemented in R, using the same methods for sparse matrices, matrix solves, and other tasks. The results were obtained using a MacBook Pro Laptop with an M3 Max processor and 128Gb of memory, without using any parallel computations to make the comparison as fair as possible.
Table 1: Asymptotic costs for different methods. For nnGP, $m$ is the number of neighbors and for the low rank methods, $m$ is the rank. For the state-space method, the “preconditioned” approximation of Karvonen and Särkkä (2016) is used in combination with our Markov approach, see Remark 11. All costs are in “Big O” assuming $n$ is much larger than $\alpha$ and $m$. For FEM, $N$ is the number of mesh nodes.
@@
state-space (Karvonen and Särkkä, 2016) $n(m+\lfloor\alpha\rfloor)\lceil\alpha\rceil^{4}$ $n(m+\lfloor\alpha\rfloor)\lceil\alpha\rceil^{2}$ $n(m+\lfloor\alpha\rfloor)^{3}\lceil\alpha\rceil^{3}$
@@
Fourier (Rahimi and Recht, 2007) $nm$ $nm$ $nm^{2}$
@@ 4.1 Setup for the first comparison
We first consider the case of a dataset with 5000 observations, ${\boldsymbol{\mathrm{y}}}=[y_{1},y_{2},\dotsc,y_{n}]$, with the observation locations being evenly spaced over the interval $I=[0,50]$. Each observation is generated as $y_{i}\mid u(\cdot)\sim\mathcal{N}(u(t_{i}),\sigma_{e}^{2}),$ where $t_{i}$ are the observation locations, and $u$ is a centered Gaussian process with the Matérn covariance function given by (1). We set $\sigma=1$ (as it is merely a scaling parameter) and explore two different noise levels: $\sigma_{e}=0.1$ and $\sigma_{e}=\sqrt{0.1}$. Additionally, we vary the smoothness parameter $\nu$ over the interval $(0,2.5)$. For each value of $\nu$, we choose $\kappa=\sqrt{2\nu}$, ensuring that the practical correlation range, $\rho=\nicefrac{{\sqrt{8\nu}}}{{\kappa}}$, remains fixed at 2.
This choice of range, along with the interval length, ensures numerical stability across all methods compared. Specifically, a practical correlation range of 2 was the largest range for which all methods, including the most sensitive (nnGP), remained stable. While a practical correlation range of 2 on a domain of size 50 may seem small, stability and accuracy of the predictions depend on the number of observations within the correlation range at any given point, rather than the size of the domain itself. Other ranges, and other numbers of observation and prediction locations, are explored in the accompanying Shiny app. For example, one scenario considers $10000$ evenly spaced prediction locations and $5000$ observation locations randomly selected without replacement from the prediction locations.
The goal is to compute the posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ with elements $({\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}})_{j}=\mathbb{E}(u(p_{j})|{\boldsymbol{\mathrm{y}}})$, and the posterior standard deviation ${\boldsymbol{\mathrm{\sigma}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ with elements $({\boldsymbol{\mathrm{\sigma}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}})_{j}=\sqrt{\operatorname{Var}(u(p_{j})|{\boldsymbol{\mathrm{y}}})}$, where $p_{j}$ are the prediction locations. To make the comparison simple, we initially choose $p_{j}=t_{j}$ evenly spaced in the interval. Because the locations are evenly spaced, we could also include the Toeplitz method of Ling and Lysy (2022) in the comparison. However, we do not include that here as we only consider generally applicable methods.
@@ 4.2 Calibration of the computational costs
The asymptotic costs of the different methods are summarized in Table 1. However, we calibrate the methods to have the same total runtime for assembling all matrices (the construction cost) and computing the posterior mean (the prediction cost). This ensures fair comparisons based on actual performance rather than theoretical complexity, which can be affected by the constants in the cost expression and the size of the study. Specifically, for a given set of parameters $(\kappa,\sigma,\nu)$, and a fixed value of $m$ for the proposed method, we calibrate the values of $m$ for the other methods to ensure that the total computation time is the same. The total prediction times were averaged over 100 samples to obtain the calibration. The final calibration results are shown in Table 2. We now provide some details regarding this calibration procedure.
@@
As described in Remark 11, the state-space method provides an alternative Markov representation for which we could use the same computational methods as for the proposed method. Its value of $m$ was therefore chosen according to Table 1 as $m-\lfloor\alpha\rfloor$.
To minimize boundary effects of the FEM method, we extended the original domain $[0,50]$ to $[-4\rho,50+4\rho]$, where $\rho$ is the practical correlation range. This extension ensures accurate approximations of the Matérn covariance at the target locations. We refer the reader to Khristenko et al. (2019, Theorem 3.2) and Bolin et al. (2024b, Proposition 2) for a theoretical justification of this commonly used procedure. Because the FEM method uses the same type of rational approximation as the propose method, we fixed the value of $m$ for the FEM method to be equal to $m$ for the proposed method. The calibration was then instead performed on the finite element mesh. Specifically, a mesh with $N=kn+500-(k+3)$ locations in the extended domain used, where $500$ locations were in the extensions and $kn$ locations in the interior $[0,50]$, and the $-(k+3)$ term appears to ensure that the regular mesh contains the observation locations. These $kn$ locations were chosen equally spaced to include the observation locations and $k\in\mathbb{N}$ was calibrated to make the total computational cost match that of the proposed method for $\nu<1.5$, and for $2.51.5<\nu<$ 2.5 it was chosen as the largest values that yielded a stable prediction, as the value which matched the computational cost yielded unstable predictions. The resulting values of $N$ are shown in Table 3.
4.3 Covariance errors
@@
$\|r_{\alpha}-\hat{r}_{\alpha}\|_{L_{2}(I\times I)}^{2}\approx\frac{1}{n^{2}}\sum_{i,j=1}^{n}(r(t_{i},t_{j})-\hat{r}(t_{i},t_{j}))^{2},\,\,\,\|r-\hat{r}\|_{C(I\times I)}\approx\max_{i,j}|r(t_{i},t_{j})-\hat{r}(t_{i},t_{j})|,$
where $r_{\alpha}$ denotes the Matérn covariance given in (5), $I=[0,50]$ is the domain we are considering, and $\hat{r}_{\alpha}$ represents the approximation obtained using the methods under consideration.
@@
Figure 3 illustrates the resulting $L_{2}$ and supremum norm errors in the covariance approximation for the different methods when $n=5000$. The left panel illustrates the error for our method with varying choices of $m$. Notably, the covariance error decreases rapidly to zero as $m$ increases, with the rate of decrease becoming faster for larger values of $\nu$. Moreover, one can observe that, especially for $\nu$ away from zero, each increase in $m$ leads to improvements on the scale of orders of magnitude. In the remaining two panels, we show the error for $m=3$ and $m=5$ and compare these errors to the error from the competing methods, calibrated to have the same cost of prediction as described above. We can note that the competing methods are less accurate and in fact, one can observe that the nnGP method becomes slightly unstable for large values of smoothness parameter $\nu$, whereas the FEM method becomes very unstable due to the calibrated mesh being very fine. It is also noteworthy that the mesh is fine enough that when the method is stable (which mean $\nu<0.9$ for $m=3$ and $\nu<0.5$ for $m=5$), we can only see the rational approximation error, as both FEM and rational methods overlap. This shows that the FEM method is an excellent choice when stable, but that the rational is better because it is more stable and does not require choosing a mesh or a boundary extension. The PCA method is more accurate for large values of $\nu$, but one should recall that it is a theoretical lower bound for low rank methods, and not a practically competitive method, as it has an $\mathcal{O}(n^{3})$ construction cost unless the eigenvectors are known explicitly. Any practically useful low rank method, such as the process convolution approach (Higdon, 2002) or fixed rank kriging (Cressie and Johannesson, 2008), will have larger errors, as can clearly be seen when considering the errors of the random Fourier features method. Additionally, it is important to note that for the values $\nu=0.5,1.5,2.5$ within the considered interval $(0,3)$, the proposed method is exact. Further, Figure 3 highlights an important observation regarding the theoretical bounds and numerical results: The numerical errors behave similar to the (approximate) theoretical bounds in Figure 2, where the error is monotonic in $m$ but not monotonic in $\{\alpha\}$.
@@
Figure 4: Prediction errors, measured as mean squared errors of the posterior means and posterior standard deviations, evaluated on the interval $I=[0,50]$ for $\sigma_{e}=0.1$, $\rho=2$, various values of $\nu$, and different orders of approximation.
4.4 Prediction errors
To compute the prediction error, we first compute the true posterior mean ${\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$, the posterior mean $\hat{{\boldsymbol{\mathrm{\mu}}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ under each approximate model and the corresponding $L_{2}$ errors $\|{\boldsymbol{\mathrm{\mu}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}-\hat{{\boldsymbol{\mathrm{\mu}}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}\|_{L_{2}}$, where $\|{\boldsymbol{\mathrm{v}}}\|_{L_{2}}^{2}=\frac{1}{n}\sum_{i=1}^{n}{\boldsymbol{\mathrm{v}}}_{i}^{2}$ for ${\boldsymbol{\mathrm{v}}}\in\mathbb{R}^{n}$. Similarly, we also compute the true posterior standard deviation ${\boldsymbol{\mathrm{\sigma}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ and the posterior standard deviation $\hat{{\boldsymbol{\mathrm{\sigma}}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}$ under each approximate model and the corresponding $L_{2}$ error $\|{\boldsymbol{\mathrm{\sigma}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}-\hat{{\boldsymbol{\mathrm{\sigma}}}}_{{\boldsymbol{\mathrm{u}}}|{\boldsymbol{\mathrm{y}}}}\|_{L_{2}}$.
Figure 5: Prediction errors, measured as mean squared errors of the posterior means and posterior standard deviations, evaluated on the interval $I=[0,50]$ for $\sigma_{e}=\sqrt{0.1}$, $\rho=2$, various values of $\nu$, and different orders of approximation.
The prediction errors under the noise levels $\sigma_{e}=0.1$ and $\sigma_{e}=\sqrt{0.1}$ are shown in Figure 4 and Figure 5, respectively. The left columns of these figures show that the $L_{2}$ error of both the posterior means and posterior standard deviations decrease consistently and rapidly to zero for the rational approximation as $m$ increases. The other two columns of the figures compare the errors of the different methods when calibrated to have the same cost as the rational approximation method with $m=3$ and $m=5$. Yhese results show that state-space method and the Fourier method are not competitive. The PCA method performs poorly for the posterior standard deviations, and also for the posterior mean unless $\nu$ is large. Additionally, using PCA at this cost requires prior knowledge of the eigenvectors, so it is typically not a feasible option. Further, the tapering method also performs poorly except for very small values of $\nu$ (essentially $\nu<0.25$ where it is not calibrated). This leaves three methods which are competitive: the rational approximation, the FEM method and nnGP.
The accuracy of the FEM method is essentially equal to that of the rational approximation for $\nu\leq 0.5$, whereas it is slightly worse for $\nu>0.5$. The nnGP method has the highest accuracy for $\nu<1$ and $\sigma_{e}=0.1$, but one should recall that the method is not calibrated for $\nu<0.5$ so the comparison is not completely fair. For $\nu>1$ and $\sigma_{e}=0.1$, the two method have similar performance but the rational approximation is slightly better for the posterior mean. For $\sigma_{e}=\sqrt{0.1}$, the rational approximation outperforms nnGP and all other methods if $\nu>1$, whereas for $\nu\in(0.5,1)$, the rational approximation has the best accuracy for the posterior mean but nnGP has the best accuracy for the standard deviation.
Further comparisons for other choices of the number of prediction locations $n$, the number of observation locations $n_{obs}$, and for different practical correlation ranges can be found in the accompanying Shiny application. The general conclusions for these choices align closely with those presented above. The main difference is that, for prediction tasks with a very small practical correlation range and an equal number of prediction and observation locations, nnGP tends to perform the best. This scenario is particularly favorable for nnGP, as all locations coincide with support points. Additionally, when $n>n_{obs}$, the rational approximation outperforms all methods, except for $\nu>1.5$, where nnGP achieves the best performance in prediction tasks.